How many grams of HCL are consumed by the reaction of 2.50 moles of magnesium?


1
Expert's answer
2021-01-27T04:40:38-0500

Reaction between Mg and HCl

Mg + 2HCl = MgCl2 + H2

Molar mass of HCl = (35.5 +1) = 36.5 gmol-1

So 1 mol Mg consume 2 mol HCl =( 2×36.5) g HCl

So 2.5 mol Mg consume = (2×36.5×2.5)

= 182.5 g HCl